OpenStudy (tia_sweetheart17):

factor out the gcf form the polynomial. 14x^3-4x^2+6x

OpenStudy (tia_sweetheart17):

@Callisto help me please

OpenStudy (solomonzelman):

\[14=2 \times 7\]\[4=2 \times 2\]\[6=2 \times 3\]what is the gcf?

OpenStudy (tia_sweetheart17):

2x

OpenStudy (solomonzelman):

Yes! Good Job!